# Multi-Robot Coordination Induced in Hazardous Environments through an Adversarial Graph-Traversal Game

### v1

@@ III-C Bounds for the Outcome
Here we provide security strategies for both players which allow us to bound the expected game outcome $J(s,\pi_{\text{blue}},\pi_{\text{red}})$, for $s\in\mathcal{S}$. While we may not know how to find the value or the Nash equilibrium policies theoretically, and approximating them numerically may be computationally intensive, these bounds will be relatively simple to compute and can be achieved by simple naive strategies.
@@ III-C1 Blue Player’s Security Strategy
This conservative bound assumes that the red player knows blue player’s actions, that it has complete freedom in choosing the next graph, and that the blue player will not use any information about ${G}(t)$ in the future. However, the strategy allows the blue team to reach the goal with a guaranteed bound on the cost.
@@ III-C2 Red Player’s Security Strategy
Under the condition (III-B), for $M=1$, if the red player follows the policy $\hat{\pi}_{\text{red}}$ defined by (III-C2), then
@@
which is the minimum discounted distance to the goal on this graph, and is the result of the blue player’s best response to this policy $\pi_{\text{red}}^{\prime}$. If condition (III-B) is satisfied, then it can be shown that shortest paths to the goal do not contain repeated nodes. Therefore, we can upper bound the number of edges traversed on such a path by $N-1$. This allows us to claim
@@
This conservative bound assumes that the red player can only change the graph at the current time step and it assumes that the blue player will choose its best response actions. However, this provides a relatively simple to compute strategy for the red player with a guaranteed bound on its payoff.
@@ III-D Game Solution
Although the problem motivation requires the blue robots to reach the goal, the introduction of the discount factor $\gamma<1$ means that the blue player can achieve finite, and potentially lower costs without doing so. In such cases, solutions to the game are not necessarily solutions to the original problem. Therefore, here we provide a condition to guarantee that, if both players use Nash equilibrium strategies $\pi_{\text{blue}}^{*}$ and $\pi_{\text{red}}^{*}$, the blue player will reach the goal.
@@
For $M=1$, given an initial condition $S(0)\in\mathcal{S}$, under the Markov chain with state transitions given by (II) with actions chosen according to $\pi_{\text{blue}}^{*},\pi_{\text{red}}^{*}$ satisfying (II), if $\gamma$ satisfies (III-B), then $\lim_{t\rightarrow\infty}\text{Pr}\left(p(t)=N\right)=1$.
@@
We present only a sketch of the proof for reasons of space. If the policies of both players are fixed to their Nash equilibrium values, the game reduces to a Markov chain. Therefore, we simply need to show that any state $s=(p,k,\alpha)\in\mathcal{S}$ where $p=N$ is an absorbing state and any other states are not absorbing for this Markov chain [19]. Then, if we can show that the Markov chain is absorbing, we have our desired result.
@@
Because we have an upper bound for the value $\overline{V}(s)$, if we can guarantee that $\overline{V}(s)<\frac{1}{1-\gamma}C_{\text{min}}$, then we know that there is a positive probability that, under Nash equilibrium policies, the blue robot reaches the goal at some future time step from state $s$. Noting that $\overline{V}(s)\leq d_{\text{max}}$, from the condition (III-B), we know that $V(s)\leq\overline{V}(s)<\frac{1}{1-\gamma}C_{\text{min}}$ for all $s\in\mathcal{S}$, so there must be a strictly positive probability that the blue robot reaches the goal from every state $s$. This also guarantees that no state with $p\neq N$ is absorbing.
@@
Therefore, by choosing $\gamma$ high enough, we can ensure the blue team will be incentivized to reach the goal.
@@ IV Numerical Methodology
Because we believe that solving for these equilibrium strategies $\pi_{\text{blue}}^{*},\pi_{\text{red}}^{*}$ analytically is not feasible within this work, we turn to numerical methods. We use a value iteration method based on the work of Shapley [17, 14]. First, we define a matrix valued Q function $\mathcal{Q}\in\mathbb{R}_{\geq}^{|A_{\text{red}}(s)|\times|A_{\text{blue}}(s)|}$ as a function of the state $s\in S$, where each element is
$\mathcal{Q}_{a_{r},a_{b}}(s)=C\left(s,a_{\text{blue}}\right)+\gamma V(s^{+}),$ (17)
where $a_{r}\in A_{\text{red}}(s)$ is an action of the red player, $a_{b}\in 1,2,\dots,|A_{\text{blue}}(s)|$ is an index corresponding to blue player’s action $a_{\text{blue}}$, and $s^{+}$ is the state at the next time step from (II), given $a_{\text{blue}}$ and $a_{\text{r}}$. This function gives the value of the game outcome if each player takes an arbitrary action at the current time step, but then follows the equilibrium policy ($\pi_{\text{blue}}^{*}$ or $\pi_{\text{red}}^{*}$) for all future time steps. According to Shapley’s Theorem, the equilibrium policies $\pi_{\text{blue}}^{*}(s)$, $\pi_{\text{red}}^{*}(s)$ and the game value can then be found by solving the matrix game corresponding to $\mathcal{Q}(s)$, for each state $s\in\mathcal{S}$, where the red player is the row player and maximizing and the blue player is the column player and minimizing [14]. However, because we do not know the value function, we instead use numerical methods to calculate the Q function.
@@
$\hat{\mathcal{Q}}^{\tau+1}_{a_{\text{b}},a_{\text{r}}}(s)=C\left(s,a_{\text{blue}}\right)+\gamma{\pi_{\text{red}}^{\tau}}^{T}\hat{\mathcal{Q}}^{\tau}(s^{+})\pi_{\text{blue}}^{\tau},$ (18)
for each $s\in\mathcal{S}$, $a_{b}\in 1,2,\dots,|A_{\text{blue}}(s)|$, and $a_{\text{r}}\in A_{\text{red}}(s)$, and where $s^{+}$ is found using (II) and $a_{\text{blue}}$ is the action corresponding to $a_{\text{r}}$. When this converges, $\pi_{\text{blue}}^{\tau},\pi_{\text{red}}^{\tau}$ can approximate the equilibrium strategies $\pi_{\text{blue}}^{*},\pi_{\text{red}}^{*}$.
@@ IV-A1 Dominated Blue Actions
Under certain circumstances, it is possible to determine that there are some edges that it is never beneficial for the blue robot to take as it traverses the graph. Specifically, we look for strictly dominated actions. For a given game state $S(t)\in\mathcal{S}$, let $p(t)=p_{0}\in\mathcal{V}$, and consider two actions of the blue player, $p_{1},p_{2}\in A_{\text{blue}}(S)$. A sufficient condition for action $p_{2}$ to dominate action $p_{1}$ is
@@
for all $a_{\text{r}}\in A_{\text{red}}(S)$, where $\alpha_{a_{\text{r}}}$ is the red player’s new ammo count as determined by $S$ and $a_{\text{r}}$ according to (II).
Because the blue player never benefits from using a dominated strategy, if condition (IV-A1) is satisfied for some $p_{0},k,p_{1},p_{2}$ and all $\alpha,{G}$, then we can remove edge $(p_{0},p_{1})$ from the graphs before solving for the value, without changing the results. Therefore, one could iterate over all pairs of actions for all nodes, removing the edges corresponding to dominated actions, and finally remove any nodes from which the goal node cannot be reached. This allows the game to be solved for fewer states and actions for the blue player, without changing the results.
@@ IV-A2 Sub-Game Formulation
From equations (4) and (II), we can write the value, for a given game state $S(t)=s=(p,{G},\alpha)\in\mathcal{S}$, as
@@
where $s^{+}=(p^{+},{G}^{+},\alpha^{+})=S(t+1)\in\mathcal{S}$ and $s^{\prime}=(p^{\prime},{G}^{\prime},\alpha^{\prime})\in\mathcal{S}$, and $\text{Pr}\{s(t+1)=s^{\prime}\}$ is found according to (II) when the actions are chosen according to Nash equilibrium policies, $\pi_{\text{blue}}^{*}$ and $\pi_{\text{red}}^{*}$. Now, the two possibilities are either the graph remains the same, in which case $s^{+}=(p^{+},{G},\alpha)$, or it is changed by the red player, in which case $s^{+}=(p^{+},{G}^{+},\alpha-1)$. This is important, because it implies that the value of the game at some ammo state $\alpha$ depends only on the value at lower ammo states $\alpha^{\prime}\leq\alpha$. We can take advantage of this to formulate sub-games which allow us to solve the original game.
Instead of considering the graph ${G}$ and the ammo $\alpha$ to be part of the game state, we consider them to be parameters of a sub-game. The game state of the sub-game is simply blue’s position $p$, and the sub-game ends when red takes action to change the graph. The actions of the players, the game dynamics, and the stage costs are the same as before, but now we consider a terminal cost when the red player ends the sub-game by changing the graph. The value for the sub-game defined by graph ${G}$ and ammo $\alpha$ is $V_{{G},\alpha}(p)$ for sub-game state $p\in\mathcal{V}$, and the terminal costs will be $V_{{G}(T),\alpha-1}(p_{T})$, where $p_{T}$ is blue’s position when red takes action to change the graph to ${G}(T)\neq{G}$ at time $T-1$. We can write the outcome of this sub-game, for $p(0)=p$, as
$V_{{G},\alpha}(p)=E\Big{[}$ $\sum_{t=0}^{T-1}\gamma^{t}C\left(S(t),p(t+1)\right)$
$+\gamma^{T}V_{{G}^{\prime},\alpha-1}(p(T))\Big{]}.$ (20)
Note that $V_{{G},0}(p)=d_{G}(p,N)$, allowing us to solve for the value of the full game by iteratively solving the sub-games, using the value of the sub-games with ammo $\alpha$ to solve for the value of the sub-games with ammo $\alpha+1$.
Compared to the full game, if we have $K$ different graphs with $N$ nodes and a maximum ammo of $\overline{\alpha}$, the trade-off is considering $K\overline{\alpha}$ sub-games with $N$ states each, instead of considering one game with $NK\overline{\alpha}$ states. Because the number of game states corresponds to the number of Q-matrices which must be updated and the number of auxiliary matrix games which must be solved at every iteration, solving the game as sub-games may result in less computation, as only the Q-matrices and value relevant to the currently considered sub-game must be updated at a time.
@@ V Numerical Results
Here we show the results of applying the numerical methods in Section IV. Unless otherwise specified, the discount factor was chosen to be $\gamma=1-10^{-9}\approx 1$ and the red player’s action graph $\mathcal{G}^{\text{red}}$ is complete. First, we show and discuss the results for some interesting cases, and then we perform a statistical analysis of randomly constructed graphs.
@@ V-A Illustrative Examples
Figure 4 shows an example graph where each player mixes between three options from node $1$, with the blue player mixing between remaining at that node and between taking either branch. With $s(0)=(1,{G^{1}},1)$, the blue player chooses its action according to $[\text{Pr}(p^{+}=1),\text{Pr}(p^{+}=2),\text{Pr}(p^{+}=3)]=[0.5,0.25,0.25]$, while the red player chooses its action according to $\pi_{\text{red}}^{*}(s(0))=[0.5,0.25,0.25]^{T}$. Intuitively, the blue player seems to mix between remaining at the current node and moving towards the goal when the current graph has a high immediate edge cost, while the red player is incentivized to change the graph because the current graph has lower costs closer to the goal. On the other hand, if ${G}(0)={G^{2}}$ instead, the blue player chooses its action according to $[\text{Pr}(p^{+}=1),\text{Pr}(p^{+}=2),\text{Pr}(p^{+}=3)]=[0,0.5,0.5]$, and the red player chooses its action according to $\pi_{\text{red}}^{*}(s(0))=[0,0.75,0.25]^{T}$. In this case, the players both only mix between two different actions, because one path is cheaper on one graph, and the other path is cheaper on the other graph.
Figure 4: This graph results in mixed strategies; node 4 is the goal node. The blue arrows indicate the possible actions of the blue player under the equilibrium strategy from node $1$. On graph $1$, the red player also mixes over its three options, but on graph 2 or 3, the red player only mixes between selecting graph 2 and graph 3.
@@
The graph in Figure 5 shows a case where having more robots is beneficial for the blue player. The only differences between the two graphs are that the $(5,7)$ edge weight is cheaper in Graph 1 and the $(6,7)$ edge weight is cheaper in Graph 2. Due to the longer length of these branches (compared to the graphs in 4), the red player can simply choose ${G}(t+1)={G^{1}}$ when $p(t)=4$ or ${G}(t+1)={G^{2}}$ when $p(t)=3$, forcing the blue player to take the higher edge cost. If the blue player has multiple robots, however, then they can split up in a coordinated fashion, so that at least one of them takes the cheaper path. To formalize this benefit, we define the average value per robot as $\tilde{V}_{M}(s)\triangleq V(s)/M$. When $\alpha(0)\geq 1$, for a single robot ($M=1$), $\tilde{V}_{1}(S(0))=20$ with $p(0)=1$, but for two robots, $\tilde{V}_{2}(S(0))=13$ with $p(0)=[1,1]^{T}$, ${G}(t)={G^{2}}$, and $\alpha(0)\geq 1$. This indicates that, depending on the graphs, the blue player can gain an advantage from having multiple robots which split up in a coordinated fashion.
Another type of coordinated behavior seen in Figure 5 is that the robot which could traverse fewer edges to reach the goal instead waits at node 5. If this leading robot on node 5 goes to the goal immediately, then the red player could switch the graph so that both robots must take the $16$ cost to the goal. Instead, that robot waits, which is more costly for itself, but since both robots traverse the final risky edges synchronously, at least one of them achieves the cheaper cost.
Figure 5: Here, it is better for the blue player to have multiple robots which split up to reach the goal node $7$. The blue arrows, shown on graph $1$, show the policy of a single robot starting at node 1, with the speech bubbles indicating the time of each action. The blue arrows shown on graph $2$ indicate the paths followed by two blue robots starting from node $1$. For $M=1$, and $\alpha(0)=1$, the red player will always change the graph so that the blue player will get a cost of $16$ for taking the $(6,7)$ edge, but for two robots the red player’s policy has no effect on the game outcome and exactly one robot will take the cheaper path to the goal.

### v2

@@ III-C Bounds for the Outcome
Here we provide security strategies for both players which allow us to bound the expected game outcome. These bounds will be simple to compute and can be achieved by simple strategies, compared to finding the game solution.
@@ III-C1 Blue Player’s Security Strategy
This simple strategy enables the blue team to reach the goal with a conservative bound on the cost.
@@ III-C2 Red Player’s Security Strategy
The red player assumes it can only change the graph at the current time and that it will remain constant in the future.
@@
Under the condition (6), for $M=1$, if the red player follows the policy $\hat{\pi}_{\text{red}}$ defined by (III-C2), then
@@
which is the minimum discounted distance to the goal on this graph, and is the result of the blue player’s best response to this policy $\pi_{\text{red}}^{\prime}$. If condition (6) is satisfied, then it can be shown that shortest paths to the goal do not contain repeated nodes. Therefore, we can upper bound the number of edges traversed on such a path by $N-1$. This allows us to claim
@@
This simple strategy provides the red player a conservative bound on its payoff.
@@ III-D Game Solution
Although the objective is for the robots to reach the goal, the introduction of the discount factor $\gamma<1$ means that the blue player can achieve finite, potentially lower costs without doing so. In such cases, solutions to the game are not necessarily solutions to the original problem. However, the following condition ensures that, if both players use NE strategies $\pi_{\text{blue}}^{*}$ and $\pi_{\text{red}}^{*}$, the blue player will reach the goal.
@@
For $M=1$, given an initial condition $S(0)\in\mathcal{S}$, under the Markov chain with state transitions given by (II) with actions chosen according to $\pi_{\text{blue}}^{*},\pi_{\text{red}}^{*}$ satisfying (II), if $\gamma$ satisfies (6), then $\lim_{t\rightarrow\infty}\text{Pr}\left(p(t)=N\right)=1$.
@@
We present only a sketch of the proof for reasons of space. If the policies of both players are fixed to their NE values, the game reduces to a Markov chain. Therefore, we simply need to show that any state $s=(p,k,\alpha)\in\mathcal{S}$ where $p=N$ is an absorbing state and any other states are not absorbing for this Markov chain [26]. Then, if we can show that the Markov chain is absorbing, we have our desired result.
@@
Because we have an upper bound for the value $\overline{J}\left(s\right)$, if we can guarantee that $\overline{J}\left(s\right)<\frac{1}{1-\gamma}C_{\text{min}}$, then we know that there is a positive probability that, under NE policies, the blue robot reaches the goal at some future time step from state $s$. Noting that $\overline{J}\left(s\right)\leq d_{\text{max}}$, from the condition (6), we know that $V(s)\leq\overline{J}\left(s\right)<\frac{1}{1-\gamma}C_{\text{min}}$ for all $s\in\mathcal{S}$, so there must be a strictly positive probability that the blue robot reaches the goal from every state $s$. This also guarantees that no state with $p\neq N$ is absorbing.
@@
Therefore, when $\gamma$ is high enough, the blue team is incentivized to reach the goal.
@@ IV Numerical Methodology
Because we believe that solving for these equilibrium strategies $\pi_{\text{blue}}^{*},\pi_{\text{red}}^{*}$ analytically is not feasible within this work, we turn to numerical methods. We use a value iteration method based on the work of Shapley [23, 20]. First, define the Q function $\mathcal{Q}(s)\in\mathbb{R}_{\geq}^{|A_{\text{red}}(s)|\times|A_{\text{blue}}(s)|}$ as a function of the state $s\in S$, where each element is
$\mathcal{Q}_{a_{r}a_{b}}(s)=C\left(s,a_{\text{blue}}\right)+\gamma V(s^{+}),$ (17)
where $a_{r}\in\left[|A_{\text{red}}(s)|\right]$ and $a_{b}\in\left[|A_{\text{blue}}(s)|\right]$ are action indices corresponding to action $a_{\text{red}}$ of the red player and $a_{\text{blue}}$ of the blue player, respectively, and $s^{+}$ is the state at the next time step from (II), given $a_{\text{blue}}$ and $a_{\text{red}}$. This function gives the expected game outcome if each player takes an arbitrary action at the current time step, and then follows the equilibrium policy in the future. According to Shapley’s Theorem, the equilibrium policies $\pi_{\text{blue}}^{*}(s)$, $\pi_{\text{red}}^{*}(s)$ and the game value can be found by solving the matrix game defined by $\mathcal{Q}(s)$, for each state $s\in\mathcal{S}$, where the red row player maximizes and the blue column player minimizes [20]. However, because the value function is unknown, we use numerical methods to calculate the Q function.
@@
$\hat{\mathcal{Q}}^{\tau+1}_{a_{\text{b}}a_{\text{r}}}(s)=C\left(s,a_{\text{blue}}\right)+\gamma{\pi_{\text{red}}^{\tau}}^{T}\hat{\mathcal{Q}}^{\tau}(s^{+})\pi_{\text{blue}}^{\tau},$ (18)
for each $s\in\mathcal{S}$, $a_{b}\in\left[|A_{\text{blue}}(s)|\right]$, and $a_{\text{r}}\in\left[|A_{\text{red}}(s)|\right]$, and where $s^{+}$ is found using (II) and $a_{\text{blue}}$ is the action corresponding to $a_{\text{b}}$. When this converges, $\pi_{\text{blue}}^{\tau},\pi_{\text{red}}^{\tau}$ can approximate the equilibrium strategies $\pi_{\text{blue}}^{*},\pi_{\text{red}}^{*}$.
@@ IV-A1 Dominated Blue Actions
It is possible to determine graph edges which are never beneficial for the blue robot to traverse, by looking for strictly dominated actions. For a given game state $S(t)\in\mathcal{S}$, let $p(t)=p_{0}\in\mathcal{V}$, and consider two actions of the blue player, $p_{1},p_{2}\in A_{\text{blue}}(S)$. A sufficient condition for action $p_{2}$ to dominate action $p_{1}$ is
@@
for all $a_{\text{r}}\in A_{\text{red}}(S)$, where $\alpha_{a_{\text{r}}}$ is the red player’s new ammo count as determined by $S$ and $a_{\text{r}}$ according to (II). Because the blue player never benefits from using a dominated strategy, if condition (IV-A1) is satisfied for some $p_{0},p_{1},p_{2}$ and all $\alpha,{G}(t)$, then edge $(p_{0},p_{1})$ can be removed from the graphs, without changing the results. Therefore, one could iterate over all pairs of actions for all nodes, removing the edges corresponding to dominated actions and any nodes from which the goal cannot be reached. This allows the game to be solved for fewer states and actions for the blue player, without changing the results.
@@ IV-A2 Sub-Game Formulation
From equations (4) and (II), we can write the value as
@@
for a given game state $S(t)=s=(p,{G},\alpha)\in\mathcal{S}$, where $s^{+}=(p^{+},{G}^{+},\alpha^{+})=S(t+1)\in\mathcal{S}$ and $s^{\prime}=(p^{\prime},{G}^{\prime},\alpha^{\prime})\in\mathcal{S}$ , and $\text{Pr}\{s(t+1)=s^{\prime}\}$ is found according to (II) when the actions are chosen according to NE policies. Now, either the graph remains the same, in which case $s^{+}=(p^{+},{G},\alpha)$, or it is changed by the red player, in which case $s^{+}=(p^{+},{G}^{+},\alpha-1)$. This implies that the value of the game at some ammo state $\alpha$ depends only on the value at lower ammo states $\alpha^{\prime}\leq\alpha$. We can take advantage of this to formulate sub-games which allow us to solve the original game.
Instead of considering the graph ${G}$ and the ammo $\alpha$ to be part of the game state, we consider them to be parameters of a sub-game with game state $p$. The sub-game ends when red takes action to change the graph. The actions of the players, the game dynamics, and the stage costs are the same as before, but now we consider a terminal cost when the graph changes. The value for the sub-game defined by graph ${G}$ and ammo $\alpha$ is $V_{{G},\alpha}(p)$, and the terminal costs will be $V_{{G}(T),\alpha-1}(p_{T})$, where $p_{T}$ is blue’s position when red takes action to change the graph to ${G}(T)\neq{G}$ at time $T-1$. The outcome of this sub-game, for $p(0)=p$, is
$V_{{G},\alpha}(p)=$
$E\Big{[}\sum_{t=0}^{T-1}\gamma^{t}C\left(S(t),p(t+1)\right)+\gamma^{T}V_{{G}(T),\alpha-1}(p(T))\Big{]}.$
Because $V_{{G},0}(p)=d_{G}(p,N)$, the value of the full game can be found by solving the sub-games, using the value of the sub-games with ammo $\alpha$ to solve for the value of the sub-games with ammo $\alpha+1$. Compared to the full game with $NK\overline{\alpha}$ states, for $K$ graphs of $N$ nodes and $\overline{\alpha}$ ammo counts, the trade-off is considering $K\overline{\alpha}$ sub-games with $N$ states each. The latter may be more computationally efficient, because the number of game states is the number of Q-matrices which must be updated and the number of matrix games which must be solved at every iteration.
@@ V Numerical Results
Here we show the results of applying the numerical methods in Section IV. The discount factor was chosen to be $\gamma=1-10^{-9}\approx 1$ and the red player’s action graph $\mathcal{G}^{\text{red}}$ is complete. First, we examine results for some examples, and then we perform a statistical analysis of randomly constructed graphs.
@@ V-A Illustrative Examples
Figure 4 shows an example graph where each player mixes between three options from node $1$, with the blue player mixing between remaining at that node and between taking either branch. With $s(0)=(1,{G^{1}},1)$, the blue player chooses its action according to $[\text{Pr}(p^{+}=1),\text{Pr}(p^{+}=2),\text{Pr}(p^{+}=3)]=[0.5,0.25,0.25]$, while the red player chooses its action according to $\pi_{\text{red}}^{*}(s(0))=[0.5,0.25,0.25]^{T}$. The blue player seems to mix between remaining at the current node and moving towards the goal when the current graph has a high immediate edge cost, while the red player is incentivized to change the graph because the current graph has lower costs closer to the goal. On the other hand, if ${G}(0)={G^{2}}$ instead, the blue player chooses its action according to $[\text{Pr}(p^{+}=1),\text{Pr}(p^{+}=2),\text{Pr}(p^{+}=3)]=[0,0.5,0.5]$, and the red player chooses its action according to $\pi_{\text{red}}^{*}(s(0))=[0,0.75,0.25]^{T}$. The players each only mix between two actions, so that the blue player attempts to take the cheaper path, which the red player attempts to prevent.
Figure 4: This demonstrates mixed equilibrium policies of both players. The arrows indicate the possible actions of the blue player under that policy from node $1$. On graph $1$, the red player also mixes over its three options, but on graph $2$ or $3$, it only mixes between selecting graph $2$ and graph $3$.
@@
The graph in Figure 5 shows a case where having more robots is beneficial for the blue player. Due to the longer length of these branches (compared to the graphs in Figure 4), the red player can simply choose ${G}(t+1)={G^{1}}$ when $p(t)=4$ or ${G}(t+1)={G^{2}}$ when $p(t)=3$, forcing the blue player to take the higher edge cost. On the other hand, multiple robots can split up in a coordinated fashion, so that at least one of them takes the cheaper path. To formalize this benefit, we define the average value per robot as $\tilde{V}_{M}(s)\triangleq V(s)/M$. When $\alpha(0)\geq 1$, for a single robot ($M=1$), $\tilde{V}_{1}(S(0))=20$ with $p(0)=1$, but for two robots, $\tilde{V}_{2}(S(0))=13$ with $p(0)=[1,1]^{T}$, and $\alpha(0)\geq 1$. This indicates that, depending on the graphs, the blue player can gain an advantage from having multiple robots which split up in a coordinated fashion. Another coordinated behavior seen in Figure 5 is that the robot which could traverse fewer edges to the goal instead waits at node $5$. If this leading robot instead went to the goal immediately, then the red player could switch the graph so that both robots must take the $16$ cost to the goal. Instead, waiting ensures that both robots traverse the final risky edges synchronously and exactly one of them achieves the cheaper cost.
Figure 5: Here, it is better for the blue player to have multiple robots which split up to reach the goal node $7$. The blue arrows on graph $1$ show the policy of a single robot, with the speech bubbles indicating the time of each action. The blue arrows on graph $2$ indicate the paths followed by two robots. For $M=1$, and $\alpha(0)=1$, the red player will always change the graph so that the blue player will get a cost of $16$ for taking the $(6,7)$ edge, but for two robots exactly one robot will take the cheaper path to the goal, no matter the red player’s action.